17,370 canned goods are to be packed in 24 boxes to be given to charity.how many boxes can be filled?how many items will be left over?
Answer by jim_thompson5910(35256)

17370/24 = 723.75

or

17370/24 = 723 remainder 18 (note: 0.75*24 = 18)

This means that the quotient is 723 and the remainder is 18.

How many boxes can be filled? 723

How many items will be left over? 18